AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 26 of 33  ELOG logo
  ID Date Author Subject Text Attachments
  162   Wed Nov 24 12:08:08 2021 CB + ML OGIon pumps onFound SIP power unit for D2000 #4 not functioning
correctly. Max current regardless of ion
element connected to.
 vacplot24.pngvacplot.pngCARMEtemps.png 
  161   Tue Nov 23 08:49:33 2021 CB + MLNEG activation completedFound temperatures and pressure stabilised
overnight. See attached.

12:00 Began activation of 3x UHV1400
 vacplot24.pngCARMEtemps.pngvacplot24-1.pngCARMEtemps-1.png 
  160   Mon Nov 22 08:06:01 2021 RSS CB + MLNEG activation starts08:20 Temperature of channels 1-3 and 5-6
were reduced to 120 degrees. Immediately
after, the conditioning of the NEGs was stopped
 IMG_5674.JPGCARMEtemps.pngvacplot.png 
  159   Fri Nov 19 12:03:51 2021 CB RSSConditionings NEGsIncreased conditioning temperature in D2000
#3 and 4 (behind detectors) to 150C from
130/140C. No appreciable change in temperature.
 CARMEtemps.pngvacplot.png 
  158   Thu Nov 18 13:50:51 2021 CB RSSThermocamera imagesUsed thermocamera provided by GSI beam
diagnostic division to take some pictures
of YR09. See attached.
 14x 
  157   Thu Nov 18 12:29:50 2021 CB RSSConditioning D2000 NEGsD2000 thermocouples report temperatures
around 100C (D2000 #1,2) and 120C (3,4).
1,2 are in the intermediate chamber, 3,4
 vacplot.pngCARMEtemps.png 
  156   Wed Nov 17 07:12:22 2021 RSS CB + ML (remotely)CARME baking with tape continues 17 NovCARME baking with heating tape continues

Left overnight at -

Ch1-3: 140 degrees
 7x 
  155   Tue Nov 16 12:05:13 2021 CB RSSCARME baking with tape continuesCARME baking with heating tape continues

Left overnight at 100C. Turned
up this morning at 9:00 to 120C.
 CARMEtemps.png20211116_132048.jpg20211116_155138.jpg20211116_155857.jpg 
  154   Fri Nov 12 08:54:11 2021 RSS ML + CB (remotely) NK AK NPFirst CARME baking attempt07:50 Small leaks (near the motors and
upper valve) were detected in the tent which
were covered with Al foil.
 Capture.PNG200W_heaters.svg40W_heater.svg 
  153   Thu Nov 11 14:01:00 2021 AK, MLHeater channel layoutsSchematic layout of heaters and channel
assignment attached. Table kindly provided
by AK. I hope the conversion from excel to
 2020.11.05_Cryring-YR09-baking-layout.pdf 
  151   Thu Nov 11 11:44:58 2021 RSS + ML OG + CB (remotely)CARME baking startedWith the help of OG, the baking tent was
built. MG started baking around 11:30 am
with a rate of 0.2 K/minute.
 IMG_5535.JPGIMG_5537.JPGIMG_5531.JPGIMG_5534.JPGIMG_5539.JPG 
  150   Wed Nov 10 18:38:16 2021 RSS TD + ML OG AK NKCARME wrapped in heating tapeHeating and electronic experts were in
CRYRING and were unhappy with the damaged
wire. Heaters were opened by OG, but no damage
 IMG_5514(1).JPGIMG_5517(1).JPGIMG_5518.JPGIMG_5520.JPGIMG_5521(1).JPG 
  149   Wed Nov 10 18:24:12 2021 RSSThermocouple labelling1. Detector

2. Chamber back

3. Strain relief
 labeling.JPG.svgLabelling.svgIMG_5512.JPG 
  147   Wed Nov 10 14:41:12 2021 TD RSS + CB (remotely)Wrapping CARME in heating tapeML reports that baking of the upstream section
(gas jet interaction chamber plus e-target)
has commenced - initially to 50 deg C to
  
  146   Tue Nov 9 15:03:29 2021 RSS + OGBurnt cables issuesUpper part of the frame was mounted. In
order to do so, the valve was closed, turbo
wire and hose were disconnected and then
 7x 
  145   Mon Nov 8 17:19:02 2021 RSS + OGNeg and ion cables connectedNeg and ion pump cables were labeled and
connected.

Kapton tape wrapped around the
 IMG_5480.JPGIMG_5483.JPGIMG_5488.JPG 
  144   Sat Nov 6 14:21:10 2021 CB RSSPreparation for bakeout tent mounting + to do for next weekConnected pneumatic hoses prepared by Gleb.
Opened flow to pressure reducer set at 2
bar. Ok. Attempted to operate pneumatic motors.
  
  143   Fri Nov 5 12:27:11 2021 CB RSS + DR NP ML USFirst pump down of YR09Tightened all flanges.

Added Pirani gauge and third IE514
to interaction chamber.
 20211105_131748.jpg20211105_135918.jpg20211105_135929.jpg20211105_140451.jpg20211105_150616.jpg 
  142   Wed Nov 3 16:39:01 2021 CB RSS + NP DR NK MLInteraction chamber lifted in placePrepared to move interaction chamber in
place. Ok CARME side, flange has M10 bolts.
On interaction chamber side i.e. bellows
 20211103_161553.jpg20211103_161813.jpg20211103_175457.jpg 
  141   Wed Nov 3 11:18:11 2021 CB RSS + DR NK MLLeak closedFound CARME at 1.0E-6 mbar. Beginning leak
test.

 
 20211103_110244.jpg20211103_110414.jpg 
ELOG V3.1.4-unknown